HUBBLE SPACE TELESCOPE DAILY REPORT #2531 PERIOD COVERED: 0000Z (UTC) 01/03/00 - 0000Z (UTC) 01/04/00 Daily Status Report as of 004/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8491 (Transition to hold, Decon and Cool Down Procedure) This program covers the transition from Protect Safe Mode to Hold Mode, then performs a Protect Decontamination. This is followed by a series of UV monitors and focus checks at -43C, then a special decon to take the instrument to -88C. The program completed nominally. 1.2 Completed Three Sets of STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.3 Completed STIS/CCD 8502 (CCD Functional for SMOV3a) The Space Telescope Imaging Spectrograph (CCD) was used to measure the baseline performance and commandability of the CCD subsystem. The proposal completed normally. 1.4 Completed Five Sets of WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 1.5 Completed STIS/MA1 8510 (FUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA1) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.6 Completed Two Sets of STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.7 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.8 Completed Four Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. The observations completed with no reported problems. 1.9 Completed FOC/96 7923 (FOC F/96 Turn-on) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 8497 (Point Spread Function Verification) The WF/PC-2 was used to verify the WFPC2 PSF after SM3A. There were no reported problems. 1.11 Completed Two Sets of STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.12 Completed STIS/CCD 8530 (SMOV3a STIS BEA Science) The Space Telescope Imaging Spectrograph (CCD) was used to obtain SMOV3a Bright Earth Avoidance time period for science exposures with the STIS CCD. Various kinds of targets are observed. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 12 Successful: 12 Scheduled Re-acquisitions: 2 Successful: 2 2.2 FHST Updates: Scheduled: 34 Successful: 34 2.3 Operations Notes: As documented in HSTAR 7458, the NICMOS instrument suspended at 003/114404Z while transitioning from boot to operate. A NICMOS 104 status buffer message was generated. A TTR was generated for a period of no ODM/GCMR capability which began at 003/1917Z. No data loss resulted. Per ROP SR-3, the SSR science playback pointer was set at 003/1955Z. The STIS MCE-2 reset at 004/065924Z while the low voltage was on and while outside any SAA interval. Using ROP NS-12, the STIS flight software error counter was reset at 004/0837Z. MAMA-2 will be recovered at 004/1950Z, the time of the next SMS-commanded high voltage-on commanding. HSTAR 7459 describes an approximate 10 percent increase in STIS CCD read noise relative to pre-SM3A values. This was discerned from data collected at 361/200004Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW